### v1

# Efficient Graph Coloring with Neural Networks: A Physics-Inspired Approach for Large Graphs

Abstract
The graph coloring problem is an optimization problem involving the assignment of one of q colors to each vertex of a graph such that no two adjacent vertices share the same color. This problem is NP-hard and arises in various practical applications. In this work, we present a novel algorithm that leverages graph neural networks to tackle the problem efficiently, particularly for large graphs. We propose a physics-inspired approach that leverages tools used in statistical mechanics to improve the training and performances of the algorithm. The scaling of our method is evaluated for different connectivities and graph sizes. Finally, we demonstrate the effectiveness of our method on a dataset of Erdős–Rényi graphs, showing its applicability also in hard-to-solve connectivity regions, where traditional methods struggle.
Keywords Graph coloring  $\cdot$ Physics-inspired Machine Learning  $\cdot$ Graph neural networks  $\cdot$ Statistical Physics  $\cdot$ Optimization Problem  $\cdot$ Potts model
1 Introduction
The graph coloring problem (GCP) is a constrained optimization problem consisting of assigning one of $q$ colors to each vertex of an undirected graph, so that there are no adjacent vertices having the same color. The problem of deciding whether a graph can be fairly colored with $q$ colors is NP-complete Karp (2010), while the optimization of a graph’s coloring to minimize conflicts is NP-hard Garey and Johnson (1979). The GCP arises in many practical scenarios, as it can applied to solve problems such as scheduling Lewis (2021), timetabling Ahmed (2012) and even solving Sudoku puzzles Tosuni (2015), moreover it can be reduced in polynomial time to all other NP-hard problems Karp (2010) as the travelling salesman problem, the clique problem or k-SAT Garey and Johnson (1979). It is widely believed that such problems cannot be solved in polynomial time owing to the $P\neq NP$ conjecture Garey and Johnson (1979) Fortnow (2009), making the GCP both difficult and useful to solve. Many current algorithms for solving the GCP rely on the local search of a solution in the configuration space Galinier and Hertz (2006) as with Monte Carlo Markov Chain (MCMC) methods like Simulated Annealing Chams et al. (1987) Johnson et al. (1991) Angelini and Ricci-Tersenghi (2023), TabuCol Hertz and Werra (1987), evolutionary algorithms Galinier and Hao (1999); Malaguti et al. (2008); Mostafaie et al. (2020). These methods, however, suffer from the exponential scaling in time, becoming inapplicable for large instances.
Other heuristic methods, like Greedy coloring Matula and Beck (1983) or DSatur Brélaz (1979), avoid the problem of time scaling but provide sub-optimal solutions. The aim of this work is presenting an algorithm that uses graph neural networks (GNNs) to search for solutions to the GCP. This paper is structured as follows:
In section 2, we provide a high-level overview of existing methods based on GNNs. In section 3 we address the GCP from a statistical mechanics point of view, highlighting its connections with the Potts model. We describe the two types of graphs used in this work, namely Erdős–Rényi and planted graphs, describing the procedure for generating them. In section 4 we detail the dataset used in this work, we present a novel architecture based on GNNs, describing the procedure used for training and coloring. In section 5 we presents the performances of our method in terms of energy and scaling at different connectivities, comparing it to simulated annealing. We show how our method is preferable to simulated annealing in a wide region of connectivities and number of nodes, in terms of speed, scalability and final energy.
2 Related Work
There is a vast and diverse list of algorithms used to solve NP-Hard problems. In this section, we focus on algorithms that primarily use Graph Neural Networks (GNNs) for solving the GCP.
In Huang et al. (2019), improved heuristics for graph coloring are sought using reinforcement learning, leveraging a deep neural network architecture that has access to the entire graph structure. In Lemos et al. (2019), a GNN is used to determine whether a graph can be colored using q colors or not. Additionally, it is shown how node embeddings can be used to assign coloring to the graph. In Li et al. (2022), the performances of aggregation-combine GNNs are studied. The work highlights the features that limit the expressive power of these models in solving node-classification problems under strong heterophily (including the GCP) and proposes solutions to improve performances.
In Schuetz et al. (2022), the authors use the architecture of a physics-inspired GNN (PI-GNN) for graph coloring. Specifically, the neural network is optimized in order to minimize the Potts energy of the given graph. This approach has been adopted in several subsequent works. In Wang et al. (2023), an alternative to PI-GNN is proposed, introducing negative-message-passing, achieving numerical improvements. Additionally, an entropic term is introduced in the loss function to accelerate convergence during the model training. Finally, in Zhang et al. (2024), a graph isomorphism network model, using a physics-inspired approach, is proposed. The output of the neural network, is post-processed using TabuCol to reduce the number of conflicts.
3 Background
In this section we formally introduce the graph coloring problem and its connection to the Potts model, furthermore the graphs ensembles used in this work are presented and the algorithm used to generate them is described.
3.1 Graph coloring
The GCP is a very well-known constraint satisfaction problem (CSP) easily understandable yet difficult to solve. Formally, we consider an undirected graph ${\mathcal{G}=(\mathcal{V},\mathcal{E})}$ where ${\mathcal{V}=\{1,...,N\}}$ is the set of vertices and =${\mathcal{E}=\{(i,j)|i,j\in\mathcal{V}\}}$ is the set of edges. The objective is to assign an integer variable (or color) ${s_{\nu}\in\{1,...,q\}}$ to each node $\nu\in\mathcal{V}$, such that nodes sharing the same edge have different colors, namely
$s_{i}\neq s_{j},\ \forall(i,j)\in\mathcal{E}$ (1)
If a graph admits such a configuration $\mathbf{s}$, the graph is said $q$-colorable. The smallest value of q for which the graph is colorable is the chromatic number $\chi(\mathcal{G})$ of the graph.
3.2 Potts model
Searching for configurations of nodes that satisfy graph coloring is equivalent to studying the energy minima of the anti-ferromagnetic Potts model Wu (1982). The Potts model is the generalization of an Ising model, where the spins can take $q$ different values. The Hamiltonian of the model is given by:
$H_{Potts}(\mathbf{s})=-\sum_{(i,j)\in\mathcal{E}}J_{ij}\delta_{s_{i},s_{j}}$ (2)
It is easy to verify that the energy value in the anti-ferromagnetic case ($J_{ij}=-1\ \ \forall(i,j)\in\mathcal{E}$), corresponds to the number of conflicting edges in the graph. In the thermodynamic limit the nodes configurations $\mathbf{s}$ for a given graph $\mathcal{G}$ in equilibrium at temperature $T$ follows the Gibbs-Boltzmann probability distribution:
$P_{\mathcal{G}}(\mathbf{s})=\frac{e^{-\beta H(\mathbf{s})}}{\mathcal{Z}}$ (3)
Where $\beta=\frac{1}{K_{B}T}$, $K_{B}$ is the Boltzmann constant and we have defined the partition function:
$\mathcal{Z}=\sum_{\{\mathbf{s}\}}e^{-\beta H(\mathbf{s})}$ (4)
In the limit $\beta\rightarrow{\infty}$, the measure is dominated by the minimum energy configurations. Therefore, the measure is the uniform distribution over the configurations that satisfy:
$H(\mathbf{s})=\min_{\mathbf{s}}H(\mathbf{s})$ (5)
Thus sampling from the distribution is equivalent to solving the GCP Mezard and Montanari (2009). This principle is behind the working of some graph coloring algorithms as simulated annealing Kirkpatrick et al. (1983) and belief propagation Pearl (2022).
It is well known that random graphs such as Random Regular Graphs (RRG) Bollobás and Bollobás (1998) and Erdős–Rényi Graphs (ERG) ERDdS and R&wi (1959) are characterized by multiple phase transitions Zdeborová and Krzakala (2007). In Fig. 1 we report a visual representation of the solution space in each phase. The success of sampling and the time required to perform it strongly depend on the average connectivity of the graph under consideration and the number of colors $q$. In particular, for sufficiently small connectivity values $c<c_{d}$, the solutions to the problem form a single connected cluster, namely it is possible to transition from one solution to another by modifying a sub-extensive number of node variables. Additionally, in this phase, finding solutions is particularly easy regardless of the algorithm used. In the $q>3$ case, for mean connectivity values $c_{d}<c<c_{c}$, there is a dynamic phase transition where an exponentially large number of clusters appears, generally slowing down the dynamics of algorithms for solving the GCP. In the region $c_{c}<c<c_{s}$, the number of clusters of zero-energy states becomes sub-exponential. In this region, despite the existence of solutions, finding them using algorithms that scale polynomially with problem size is particularly challenging. Finally, in the region with high connectivity values $c>c_{s}$, no graphs have a solution that satisfies perfect coloring. We highlight that this holds only in the thermodynamic limit, while at finite size, it remains possible to generate ERG with a solution even at high values of connectivity. The $q\leq 3$ case is less significant, as the $c_{d}$ and $c_{c}$ critical thresholds coincide, resulting in a second-order phase transition, which is generally easier to tackle for algorithms.
Figure 1: Phase transitions in the graph coloring problem from Zdeborová and Krzakala (2007).
4 Methods
In this section, we describe the model architecture and dataset. The loss function and training method are then introduced. Finally, the algorithm used for searching the solution is presented and motivated.
4.1 Model
We are interested in learning a map $f(\cdot;\mathcal{G})$ from a feature space $\mathcal{X}\in\mathbb{R}^{k\times N}$ to a color space $\{1,\dddot{},q\}^{N}$ such that:
$H_{Potts}\big{[}f(\{\mathbf{x}_{i}\}_{i\in\mathcal{V}};\mathcal{G})\big{]}=0$ (6)
We choose to parameterize this map with a GNN model, implementing the transformation $f_{\boldsymbol{\theta}}(\cdot;\mathcal{G})$. The first choice model to perform this task are GNNs, which are designed to leverage the invariance properties of graph structures, making them effective for learning representations that respect the underlying graph topology.
In recent years, a plethora of architectures have been presented in the context of GNNs Zhou et al. (2021). For our purposes, we have focused on one of the most general and expressive models, described in Battaglia et al. (2018), and already used with promising results Tang et al. (2023); Wieder et al. (2020); Wang et al. (2019) in solving various tasks. The model used in this work is based on a message passing network (MMPN) composed by $L$ layers. At each layer nodes features are updated as follows: at first for each edge an ingoing and an outgoing message are computed, (each node sends and receives a message from its neighbors) using the formula:
$\textbf{m}_{ij}^{(l)}=\phi^{(l)}_{\theta_{1}^{(l)}}\left(\mathbf{x}_{i}^{(l-1)},\mathbf{x}_{j}^{(l-1)},\mathbf{e}_{j,i}\right)$ (7)
where $\textbf{m}_{ij}^{(l)}$ represents the message sent by the $\text{j}^{\text{th}}$ node and gathered by the $\text{i}^{\text{th}}$ node in the $l^{\text{th}}$ layer, while $x_{i}^{(l-1)}$ represents the node features at the previous layer. Next, the messages are aggregated using any permutation invariant function (e.g. sum, average, max). Together with the node features from the previous step $\mathbf{x}_{i}^{(l-1)}$, the aggregated messages are used to update the node features:
$\mathbf{x}_{i}^{(l)}=\gamma^{(l)}_{\theta_{2}^{(l)}}\left(\mathbf{x}_{i}^{(l-1)},\bigoplus_{j\in\mathcal{N}(i)}\textbf{m}_{ij}^{(l)}\right)$ (8)
The function $\phi^{(k)}_{\theta_{1}^{(l)}}$ used to compute the messages and the function $\gamma^{(l)}_{\theta_{2}^{(l)}}$ needed to update the node features are parameterized by two multi layer perceptrons (MLP). After the $L^{\text{th}}$ layer, all hidden and input features are passed to a third MLP $\Gamma_{\theta_{3}}$ to compute the final output of the node:
$\mathbf{x}_{i}^{output}=\text{softmax}\Big{[}\Gamma_{\theta_{3}}(\{\mathbf{x}^{(l)}_{i}\}_{l=0,\dots,L})\Big{]}$ (9)
The softmax activation makes the output feature vector normalized. This way the $a^{\text{th}}$ component can be interpreted as a node’s probability of having the $a^{\text{th}}$ color. We denote the overall function implemented by the parametric model with $f_{\boldsymbol{\theta}}\big{(}\{\mathbf{x}_{i}^{(0)}\}_{i\in\mathcal{V}}\big{)}$, where $\boldsymbol{\theta}$ represents the set of all trainable parameters and we denote the input feature vector associated to the $i^{\text{th}}$ node with $\mathbf{x}_{i}^{(0)}$. We highlight that by construction, the hidden features $\mathbf{x}_{i}^{(l)}$ contain information about all neighbors at distance $l$ from the $\text{i}^{\text{th}}$ node. Moreover the output function $\Gamma_{\theta_{3}}$ is able to use all intermediate features at once, effectively learning to process messages gathered at different graph distances.
4.2 Dataset
We are interested in training our algorithm using both supervised and unsupervised loss terms. To do so, we need a dataset of graphs satisfying the GCP. In this work, we focused on Erdős-Rényi graphs, as there is an efficient method to generate them without hetero-chromatic edges.
4.2.1 Erdős–Rényi random graphs
An Erdős–Rényi graph $\mathcal{G}(N,M)$ is a type of random graph generated by selecting exactly $M$ edges uniformly at random from the set of all possible $\binom{N}{2}$ edges. The number of edges is connected to the graph average connectivity $c$ and number of nodes $N$ through the formula:
$M=\frac{N\cdot c}{2}$ (10)
4.2.2 Quiet planting solutions
As we have discussed in section 1 and in paragraph 3.2, finding a perfect coloring for a generic graph can be very resource-expensive. To this end, in the present paragraph, we briefly discuss a simple and effective method, proposed for the first time in Krzakala and Zdeborová (2009), that allows us to generate graphs for which a solution $\mathbf{s}^{*}$ is already known. The procedure is explained in algorithm 1.
Algorithm 1 Quiet planting algorithm
1:Input: Number of nodes $N$, number of colors $q$, number of edges $M=\frac{cN}{2}$
2:Assign each node $i\in\mathcal{V}$ a color $s_{i}\in\{1,\dots,q\}$ such that each color appears $\sim\frac{N}{q}$ times
3:Initialize an empty edge set $|\mathcal{E}|=0$
4:while $|\mathcal{E}|<M$ do
5:     Sample nodes $i,j\in\mathcal{V}$ uniformly at random
6:     if $s_{i}\neq s_{j}$ and $(i,j)\notin\mathcal{E}$ then
7:         Add edge $(i,j)$ to $\mathcal{E}$
8:     else
9:         Sample new nodes $i,j\in\mathcal{V}$ uniformly at random
10:     end if
11:end while
Notably, this algorithm can generate fairly colored graphs at any connectivity, even beyond the threshold $c>c_{s}$. It is possible to verify that these kinds of graphs are indistinguishable from graphs sampled from the Gibbs distribution when $c<c_{c}$ Zdeborová and Krzakala (2007), for this reason, this method is called quiet-planting. When $c>c_{c}$ the entropy of planted solutions becomes larger than the entropy of all other random solutions, making them distinguishable. However, the model that we will present in section 4 does not aim to sample graphs from the Gibbs distribution, but to find a solution that minimizes the energy. Therefore, the solutions obtained by quiet planting continue to be suitable for the purpose of training.
4.3 Training and loss function
In this work, we are interested in training a model that, given a colored graph with conflicting edges, produces a perfect coloring. Hence we are interested in finding a map:
$f:\{1,\dddot{},q\}^{N}\longrightarrow\{1,\dddot{},q\}^{N}$ (11)
To this end, the model has been trained with semi-supervised loss function, including an energy term and a term which measures the closeness of the output with the perfectly colored solution. The training is performed using a gradient-descent-based algorithm, therefore the loss function must be differentiable with respect to the model’s parameters. This is not possible by using the Potts energy directly, as it is not differentiable. For this reason, regarding the energy term in the loss function, we have employed a continuous version of the Potts energy, introduced in Schuetz et al. (2022):
$H_{Potts}^{c}\big{[}\{\mathbf{y}_{i}\}_{i\in\mathcal{V}}\big{]}=\sum_{(i,j)\in\mathcal{E}}\langle\textbf{y}_{i},\mathbf{y}_{j}\rangle$ (12)
where the $a^{th}$ component of $\mathbf{y}_{i}$ is the probability of the $a^{th}$ color to be assigned to the $i^{th}$ node. For this reasons both the spaces described in equation 11 are extended to $\mathbb{R}^{q\times N}$.
Regarding the supervised term in the loss function, this is inspired by a denoising autoencoder Vincent et al. (2008), which forces the model to map corrupted inputs by adding noise (thus moving them away from the solution clusters) into inputs belonging to the solution manifold. In our case, the model is trained by taking a planted solution $\{\boldsymbol{\xi}_{i}\}_{i\in\mathcal{V}}$ to which noise $\{\boldsymbol{\epsilon}_{i}\}_{i\in\mathcal{V}}\sim\mathcal{N}(\boldsymbol{0},\hat{I})$ is added
$\{\tilde{\mathbf{x}}_{i}\}_{i\in\mathcal{V}}=\sqrt{\alpha}\{\boldsymbol{\xi}_{i}\}_{i\in\mathcal{V}}+\sqrt{1-\alpha}\{\boldsymbol{\epsilon}_{i}\}_{i\in\mathcal{V}}$ (13)
and including in the loss function a term which measures the closeness of the input planted graph to the output of the model. A third contribution to the loss is given by an entropy term, measuring the average output node entropy. This term is used only during the initial steps of training to help the algorithm escape the region of parameter space minimum that results in a paramagnetic configuration $\Big{(}\big{[}f_{\boldsymbol{\theta}}(\{\textbf{x}_{i}\}_{i\in\mathcal{V}})\big{]}_{j}^{a}=\frac{1}{q},\forall j\in\mathcal{V},\forall a\in\{1,\dddot{},q\}\Big{)}$. We have empirically observed that this term doesn’t lead to a final lower energy but increases the convergence speed. The final loss function is then given by:
$L(\{\textbf{x}_{i}\}_{i\in\mathcal{V}},\{\boldsymbol{\xi}_{i}\}_{i\in\mathcal{V}};\boldsymbol{\theta})=$
$=$ $\mathbb{E}_{\{\boldsymbol{x}_{i}\}_{i\in\mathcal{V}}}\big{[}h\left[f_{\boldsymbol{\theta}}(\{\textbf{x}_{i}\}_{i\in\mathcal{V}})\right]\big{]}$
$\quad+\eta_{1}$ $\mathbb{E}_{\{\boldsymbol{x}_{i}\}_{i\in\mathcal{V}}}\big{[}S\left[f_{\boldsymbol{\theta}}(\{\textbf{x}_{i}\}_{i\in\mathcal{V}})\right]\big{]}$
$\quad-\eta_{2}$ $\mathbb{E}_{\{\boldsymbol{x}_{i}\}_{i\in\mathcal{V}}}\big{[}O\left[f_{\boldsymbol{\theta}}(\{\textbf{x}_{i}\}_{i\in\mathcal{V}}),\{\boldsymbol{\xi}_{i}\}_{i\in\mathcal{V}}\right]\big{]}$
where $\eta_{1}$ and $\eta_{2}$ are relative weights and the three contributions are respectively an energy term, namely the percentage of conflicting edges, the features entropy and the overlap with the planted graph. These terms are given by:
$\begin{cases}h\left[f_{\boldsymbol{\theta}}(\{\textbf{x}_{i}\}_{i\in\mathcal{V}})\right]=\frac{\sum_{ij}A_{ij}\langle f_{\boldsymbol{\theta}}(\boldsymbol{x}_{i}),f_{\boldsymbol{\theta}}(\boldsymbol{x}_{j})\rangle}{\sum_{ij}A_{ij}}\vspace{0.2cm}\\ S\left[f_{\boldsymbol{\theta}}(\{\textbf{x}_{i}\}_{i\in\mathcal{V}})\right]=\sum_{i\in\mathcal{V}}f_{\boldsymbol{\theta}}(\boldsymbol{x})_{i}\cdot log_{2}f_{\boldsymbol{\theta}}(\boldsymbol{x}_{i})\vspace{0.2cm}\\ O\left[f_{\boldsymbol{\theta}}(\{\textbf{x}_{i}\}_{i\in\mathcal{V}}),\boldsymbol{\xi}_{i}\right]=\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}\langle f_{\boldsymbol{\theta}}(\boldsymbol{x_{i}}),\boldsymbol{\xi}_{i}\rangle\end{cases}$
where $A_{ij}$ is the adjacency matrix and
$h\left[\{\textbf{x}_{i}\}_{i\in\mathcal{V}}\right]=\frac{1}{M}H_{Potts}^{c}\big{[}\{\mathbf{x}_{i}\}_{i\in\mathcal{V}}\big{]}$ (14)
The overlap term is necessary to break the permutation symmetry of the colors. In fact, the energy and entropy terms in the loss are invariant for color permutations. Consequently the loss landscape contains $q!$ identical regions, each corresponding to equivalent, permuted colorings. Breaking this symmetry simplifies the map that the model needs to learn, as the loss landscape is modified so that only the planted solution has the lowest loss, guiding the optimizer more efficiently.
The training phase of our model is detailed in algorithm 2.
Algorithm 2 Training algorithm (batch size = $1$)
1:for $n=1,\ldots,N_{epochs}$ do
2:     Sample: $\alpha\sim\text{Uniform}(\alpha_{\text{min}},\alpha_{\text{max}})$ [^1] In the batched case, $\alpha$ is also batched.
3:     Sample planted graph: $\{\boldsymbol{\xi}_{i}\}_{i\in\mathcal{V}}$
4:     Sample noise: $\{\boldsymbol{\epsilon}_{i}\}_{i\in\mathcal{V}}\sim\mathcal{N}(\boldsymbol{0},\hat{I})$
5:     Compute:
6:          $\{\tilde{\mathbf{x}}_{i}\}_{i\in\mathcal{V}}=\sqrt{\alpha}\{\boldsymbol{\xi}_{i}\}_{i\in\mathcal{V}}+\sqrt{1-\alpha}\{\boldsymbol{\epsilon}_{i}\}_{i\in\mathcal{V}}$
7:     Forward pass through GNN:
8:          $\{\mathbf{y}_{i}\}_{i\in\mathcal{V}}=f_{\boldsymbol{\theta}}(\{\tilde{\mathbf{x}}_{i}\}_{i\in\mathcal{V}})$
9:     Compute loss: $L(\{\textbf{x}_{i}\}_{i\in\mathcal{V}},\{\boldsymbol{\xi}_{i}\}_{i\in\mathcal{V}};\boldsymbol{\theta})$
10:     Backpropagate to compute:
11:              $\nabla_{\boldsymbol{\theta}}L(\{\textbf{x}_{i}\}_{i\in\mathcal{V}},\{\boldsymbol{\xi}_{i}\}_{i\in\mathcal{V}};\boldsymbol{\theta})$
12:     Update parameters
13:end for
Both $\alpha_{\text{max}}$ and $\alpha_{\text{min}}$ are hyperparameters of the algorithm, which have led to the best results when set respectively to 0.9 and 0.4.
4.4 Coloring
In this paragraph, we detail the coloring process using our trained model $f_{\boldsymbol{\theta}}\big{(}\cdot\big{)}$. For a sufficiently expressive model, we would expect to find the perfect coloring with a single forward pass. For our model, however, we have empirically observed that this is not the case, as the output energy is further reduced with subsequent applications of the same trained model. Additionally, we have observed that the algorithm used is characterized by fixed points that do not correspond to the solutions of the GCP. This causes the algorithm to get stuck in sub-optimal solutions. We have verified that performances are greatly improved by adding noise after each forward step, as described in the equation 13. We experimented with different choices for noise scheduling, the best performances were achieved with linearly increasing values of alpha between $\alpha_{min}=0.4$ and $\alpha_{max}=0.9$. We speculate how, during this procedure, the neural network starts by lowering the energy of the input graph. If performed for a large enough number of iterations, this procedure eventually leads the graph into the attraction basin of a zero-energy solution. The scheduling of noise from large to smaller values, has the meaning of letting the model explore a larger portion of the color space in the initial steps, while making the model collapse towards the zero-energy solution in the final steps. We have empirically observed how this procedure produces very similar results even when a colored solution is not guaranteed to exist. A more detailed discussion on how noise modifies the forward step is discussed in Appendix A. A detailed description of the coloring procedure is given in algorithm 3.
Algorithm 3 Graph coloring algorithm
1:Initialize random coloring: $\{\mathbf{x}^{(0)}_{i}\}_{i\in\mathcal{V}}$
2:Set: $\alpha_{\text{arr}}=\text{Linspace}(\alpha_{\text{min}},\alpha_{\text{max}},T)$
3:for $t=0,\ldots,T-1$ do
4:     Sample Gaussian noise: $\{\boldsymbol{\epsilon}_{i}\}_{i\in\mathcal{V}}\sim\mathcal{N}(\boldsymbol{0},\hat{I})$
5:     Compute:
6:      $\{\tilde{\mathbf{x}}^{(t)}_{i}\}=\sqrt{\alpha_{\text{arr}}[t]}\{\mathbf{x}^{(t)}_{i}\}+\sqrt{1-\alpha_{\text{arr}}[t]}\{\boldsymbol{\epsilon}_{i}\}$
7:     Forward step:
8:          $\{\mathbf{x}^{(t+1)}_{i}\}_{i\in\mathcal{V}}=f_{\boldsymbol{\theta}}(\{\tilde{\mathbf{x}}_{i}\}_{i\in\mathcal{V}})$
9:end for
5 Results
In this section we show the performances of our method in terms of conflicting edges and number of iterations at different connectivities and number of nodes, both for planted and random graphs. In the following, every data point is obtained by processing $10$ graphs extracted randomly from the test set, and reporting the mean and standard deviation.
5.1 Setup
To train our model, we have used a dataset of $1000$ planted graphs for each connectivity in $[12.5,12.6,12.7,\dddot{},14.9,15]$, for a total of $26k$ graphs, each graph in the dataset having $1000$ nodes. We have used $90\%$ of the dataset to train our model and $10\%$ for validation. We have tested our model on independently generated samples with a larger number of nodes. The results presented in the following have been obtained with a model having $5$ layers and using a latent space dimension of $32$. Furthermore we have used $q=5$ colors and an additional input feature to encode the nodes degree. We have observed that this additional feature increases the model performance with a negligible computational cost. The number of parameters of each MLP in the model is reported in table 1.
Layer Parameters Layer Parameters
$\phi_{\theta_{1}^{(1)}}^{(1)}$ 2673 $\gamma_{\theta_{2}^{(1)}}^{(1)}$ 3564
$\phi_{\theta_{1}^{(1)}}^{(2)}$ 4455 $\gamma_{\theta_{2}^{(2)}}^{(2)}$ 4455
$\phi_{\theta_{1}^{(1)}}^{(3)}$ 4455 $\gamma_{\theta_{2}^{(3)}}^{(3)}$ 4455
$\phi_{\theta_{1}^{(1)}}^{(4)}$ 4455 $\gamma_{\theta_{2}^{(4)}}^{(4)}$ 4455
$\phi_{\theta_{1}^{(1)}}^{(5)}$ 4455 $\gamma_{\theta_{2}^{(5)}}^{(5)}$ 4455
$\Gamma_{\theta_{3}}$ 6968
Total 48845
Table 1: Number of Parameters in Each Layer of our model.
We trained this model for $2000$ epochs using a batch size of $64$, $\eta_{1}=0.5$, and $\eta_{2}=0.05$, setting $\eta_{2}$ to zero after the first epoch. We have used the Adam optimizer Kingma and Ba (2014) for parameter optimization.
5.2 Comparison with simulated annealing
Here we provide a comparison of our method and simulated annealing (SA) for random and planted graphs having $10k$ nodes for different values of connectivity. As pointed out in Angelini and Ricci-Tersenghi (2023), there is no “standard benchmarks based on really hard problems". We suggest to test future proposed algorithms on instances of Erdős–Rényi graphs for connectivities ranging from $11.0$ to $15.5$, as all phase transitions for $q=5$ are explored Zdeborová and Krzakala (2007). We have chosen to compare our method against SA as it is one of the most robust and best-scaling algorithms to solve the GCP Johnson et al. (1991). Each coloring has been performed for a different number of iterations to assess the scaling of conflicting edges even when the two algorithms are executed on different hardware (respectively CPU and GPU), and a time comparison would be unfair. In figure 2 we report the results for four values of the connectivity, namely $11.5,13.0,13.5$ and $15.5$. The values were chosen in order to show the differences in each region of the phase-diagram. The comparison for remaining connectivities is reported in figure 8 in appendix B. In figure 9 in appendix C we also report the same graphs comparing execution time.
(a) Comparison at connectivity 11.5
(b) Comparison at connectivity 13.0
(c) Comparison at connectivity 13.5
(d) Comparison at connectivity 15.5
Figure 2: Comparison of the scaling of conflicting with the number of iterations
Figure 2 shows how our method consistently outperforms SA for connectivities up to $15.5$ for a number of iterations up to $102400$. For a deeper understanding of the scaling properties of our method with respect to SA, we have fitted the percentage of unsatisfied edges reached after a varying number of iterations, with respect to the number of iterations, finding that data points can be fitted very well with a power function:
$f(x)=Ax^{-B}+C$ (15)
In figure 3 we report the values of the fitted parameters for all connectivity values. The three parameters can be interpreted as follows: C represents the plateau value, namely the fitted percentage of unsatisfied edges for a large number of iterations. It is interesting to note how the fitted values for C are compatible with 00 for connectivities below or equal to $c=12.5$. This can be interpreted as the ability of both SA and our model to reach the zero-energy solution. This is compatible with the presence of the phase transition $c_{d}=12.837$, after which new solution clusters start to appear, slowing GC algorithms. We highlight how our method reaches lower plateau values than SA for all connectivities below $14.5$, while reaching compatible values for $c=14.5$ and $c=15.0$. We stress that the training was performed using connectivities uniformly distributed between $12.5$ and $15$, as detailed in paragraph 4.2. Parameter $A$ can be computed as $A=f(1)-C$, namely it represents the $y$ value of the fitted function at the first step, measured from the plateau value. Parameter $B$ represents how quickly the final energy scales with the number of iterations. In this regard, our algorithm outperforms SA for connectivities below or equal to $c=12.5$, namely it requires less iterations to reach the same final energy value. Results between the two methods are compatible for higher connectivities.
(a) Fit of A parameter for different connectivities
(b) Fit of B parameter for different connectivities
(c) Fit of C parameter for different connectivities
Figure 3: Scaling of fit parameters $A$, $B$ and $C$. $c_{c}$, $c_{d}$, $c_{r}$ and $c_{s}$ correspond to phase transitions, while $[c_{min},c_{max}]$ is the training set connectivity range.
5.3 Model scaling
By construction, our algorithm uses parallel computation to calculate messages and update all nodes simultaneously, exploiting the power of modern GPUs. This allows it to scale effectively with the size of the problem. In table 2 we report the average number of iterations per second of our algorithm, for different number of nodes and connectivities. This table shows how the number of iterations per second for our algorithm scales sub-linearly with the size of the problem, up to a size of $30k$ nodes, after which the speed decreases dramatically. On the other hand, we are interested in assessing whether our algorithm generalizes well for different number of nodes, in terms of the percentage of conflicting edges reached. For this purpose, we have processed graphs with connectivities $11.5$, $13.0$, $13.5$ and $15.5$, for $1k$, $3k$, $10k$, $30k$ and $100k$ nodes, with number of iterations distributed exponentially between $100\cdot 2^{0}$ and $100\cdot 2^{10}$. In figure 4 we report the percentage of conflicts obtained by coloring planted and random graphs for varying connectivities and number of nodes.
Figure 4: Scaling of the fraction of conflicts for different number of nodes. Planted graphs (left) and random graphs (right) are displayed side by side, for different connectivity values.
The left column of figure 4 shows how, for random graphs with $c=13.0$, $c=13.5$ and $c=15.5$, the number of conflicting edges is independent of the graph size within errors, while for $c=11.5$ smaller graphs are solved in less iterations. In the right column of figure 4, for planted graphs with $c=13.0$ and $c=13.5$, the percentage of conflicting edges is compatible for different number of nodes, while for $c=11.5$ and $c=15.5$ coloring smaller graphs results in a lower percentage of conflicts. This shows how it is possible to train our model only on a dataset of small graphs and apply it to much larger graphs maintaining satisfying results. Furthermore, for $c=13.0$ and $c=13.5$ we have shown that the percentage of conflicting edges doesn’t depend on the graph size, while for $c=11.5$ (easy to solve) and $c=15.5$ (hard to solve) smaller graphs require fewer iterations. Moreover, for $c=15.5\gg c_{c}$ the model is better at solving planted graphs rather than random ones. A possible explanation is that, for connectivities $c>c_{c}$, random and planted graphs become distinguishable, hence the model performs better on the type of graphs used for training.
6 Conclusions
In this work, we presented an algorithm that leverages physics-inspired graph neural networks to tackle the graph coloring problem. We have implemented a novel approach that uses planted solutions to introduce a supervised term in the loss function and exploited an architecture that collectively processes features at different graph distances. This architecture has been used in an iterative way, decreasing the model depth, so to avoid the common problem of over smoothing in GNNs. In this work, we have assessed the performances of our model against those of simulated annealing, as it is one of the most robust and best performing algorithms for graph coloring. Our study shows how the percentage of conflicting edges reached by our algorithm for different number of iterations, scales better than simulated annealing, on connectivities up to $c=15.5$. This region includes all phase transitions of the Potts model for $q=5$. We have shown how, up to the saturation of GPU resources, the execution time of our algorithm has a weak dependence on the size of the problem. Furthermore, for some connectivities, the algorithm performances don’t significantly change across different graph sizes. In conclusion, this work highlights how graph neural networks can be efficiently employed to tackle the graph coloring problem and scaled to very large graphs, exploiting GPUs’ ability to perform parallel computations. Future work include training larger models on a wider range of connectivities and sizes, exploring different architectures and our algorithms on real-world and other NP-hard problems.
Acknowledgments
This work was supported by PNRR MUR project PE0000013-FAIR.
Appendix A Effects of noise during coloring
Our algorithm is entirely agnostic to the way we perform coloring. We observed that a single application of our trained model doesn’t solve, in general, the graph. Additionally, as detailed in section 4.4, we observed that by repeatedly applying the same map $\mathbf{f}_{\boldsymbol{\theta}}(\cdot,\mathcal{G})$ (trained model) to an input state, the energy decreases, reaching a fixed point, namely there are solutions that satisfy $h\big{[}f_{\theta}(\{\mathbf{x}^{FP}_{i}\}_{i\in\mathcal{V}})\big{]}\approx h\big{[}\{\mathbf{x}^{FP}_{i}\}_{i\in\mathcal{V}}\big{]}$. These fixed points rarely correspond to zero-energy solutions. In this regard, we empirically observed a performance improvement when corrupting the algorithm’s output before every new step, namely when using:
$\mathbf{x}^{(t+1)}=f_{\theta}(\sqrt{\alpha}\mathbf{x}^{(t)}_{i}+\sqrt{1-\alpha}\boldsymbol{\epsilon},\mathcal{G})$ (16)
where $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\hat{I})$. In figure 5(b) the two coloring procedures are compared for different values of connectivity.
(a) Comparison at connectivity 11.5
(b) Comparison at connectivity 13.0
(c) Comparison at connectivity 13.5
(d) Comparison at connectivity 15.5
Figure 5: Percentage of conflicting edges at each iteration step during a coloring loop for different connectivities. Coloring with (without) noise is represented in red (blue). Two standard deviations are reported with a lighter coloring.
From the figure it is evident how adding adding noise during the coloring loop greatly decreases the final energy reached by our algorithm. Here we aim at providing a possible explanation to this behavior. Let’s consider zero-energy solutions $\boldsymbol{\xi}$ and fixed point solutions $\mathbf{x}^{FP}$, as defined above. We are interested in showing what happens when we apply our trained model to these solutions, after they are corrupted with noise, namely we look at:
$\begin{cases}\tilde{\boldsymbol{\xi}}_{i}(\alpha)&=f_{\theta}(\sqrt{\alpha}\boldsymbol{\xi}_{i}+\sqrt{1-\alpha}\boldsymbol{\epsilon},\mathcal{G})\\ \tilde{\mathbf{x}}^{FP}_{i}(\alpha)&=f_{\theta}(\sqrt{\alpha}\mathbf{x}^{FP}_{i}+\sqrt{1-\alpha}\boldsymbol{\epsilon},\mathcal{G})\end{cases}$ (17)
where $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\hat{I})$. We are interested in studying the behavior of the percentage of conflicts in the output configuration:
$h_{\alpha}\big{[}\{\tilde{\boldsymbol{\xi}}_{i}\}_{i\in\mathcal{V}}\big{]}=\frac{1}{M}\sum_{(i,j)\in\mathcal{E}}\langle\tilde{\boldsymbol{\xi}}_{i},\tilde{\boldsymbol{\xi}}_{j}\rangle$ (18)
In figure 6(a) we report the difference in the percentage of conflicts of a zero-energy solution $\{\boldsymbol{\xi}_{i}\}_{i\in\mathcal{V}}$ and the same solution after being corrupted with noise and processed by our trained model $\{\tilde{\boldsymbol{\xi}}_{i}\}_{i\in\mathcal{V}}$. In figure 6(b) we report the same graph for fixed point solutions $\{\mathbf{x}^{FP}_{i}\}_{i\in\mathcal{V}}$. In the case of zero-energy solution we can observe how, for sufficiently low noise levels (high $\alpha$ values), the difference in the percentage of conflicts remains very low. The noisy configurations, although differing from the original solution, continue to belong the same basin and maintain an energy statistically compatible with zero. On the other end, for fixed point solutions $\{\mathbf{x}^{FP}_{i}\}_{i\in\mathcal{V}}$, there are certain values of $\alpha$ for which the energy of the corrupted graph after the application of our model $\{\tilde{\mathbf{x}}^{FP}_{i}\}_{i\in\mathcal{V}}$ is significantly lower than the percentage of conflicts of the fixed point solution $\{\mathbf{x}^{FP}_{i}\}_{i\in\mathcal{V}}$. This suggests that in these cases, the noise allows these solutions to exit the attraction basin of the fixed point and find lower energy configurations.
(a) Average and standard deviation of $\Delta h_{\alpha}=h_{\alpha}\big{[}\{\tilde{\boldsymbol{\xi}}_{i}\}_{i\in\mathcal{V}}\big{]}-h_{\alpha}\big{[}\{\boldsymbol{\xi}_{i}\}_{i\in\mathcal{V}}\big{]}$ as a function of $\alpha$.
(b) Average and standard deviation of $\Delta h_{\alpha}=h_{\alpha}\big{[}\{{\tilde{\mathbf{x}}^{FP}}_{i}\}_{i\in\mathcal{V}}\big{]}-h_{\alpha}\big{[}\{{\mathbf{x}^{FP}_{i}}\}_{i\in\mathcal{V}}\big{]}$ as a function of $\alpha$.
Figure 6: Difference between the percentage of conflicts of a configuration before and after the application of the trained model, for $N=10k$ and $c=11.5,13.0,13.5,15.5$.
It is also interesting to study the behavior of the overlap between the original and processed graph, both for zero-energy solutions and for fixed point solutions:
$\begin{cases}O\big{[}\{\tilde{\boldsymbol{\xi}}_{i}\}_{i\in\mathcal{V}},\{\boldsymbol{\xi}_{i}\}_{i\in\mathcal{V}}\big{]}&=\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}\langle\tilde{\boldsymbol{\xi}}_{i},\boldsymbol{\xi}_{i}\rangle\\ O\big{[}\{\tilde{\textbf{x}}^{FP}_{i}\}_{i\in\mathcal{V}},\{\textbf{x}^{FP}_{i}\}_{i\in\mathcal{V}}\big{]}&=\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}\langle\tilde{\textbf{x}}^{FP}_{i},\textbf{x}^{FP}_{i}\rangle\end{cases}$ (19)
Furthermore we are interested in the difference between these two quantities, namely:
$\Delta O\big{[}\{\mathbf{x}^{FP}_{i}\}_{i\in\mathcal{V}},\{\tilde{\mathbf{x}}^{FP}_{i}\}_{i\in\mathcal{V}},\{\boldsymbol{\xi}_{i}\}_{i\in\mathcal{V}},\{\tilde{\boldsymbol{\xi}}_{i}\}_{i\in\mathcal{V}}\big{]}=\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}\langle\tilde{\mathbf{x}}^{FP}_{i},\mathbf{x}^{FP}_{i}\rangle-\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}\langle\tilde{\boldsymbol{\xi}}_{i},\boldsymbol{\xi}_{i}\rangle$ (20)
In figure 7(a) we report the overlap between a graph’s configuration and the graph obtained adding noise and applying our trained model, as a function of $\alpha$ for different values of connectivity, both for zero-energy solutions and for fixed point solutions. In figure 7(b) we report the overlap difference between the fixed point solution and zero-energy solutions for corresponding values of $\alpha$. As the figures suggests, there are values of $\alpha$ for which the zero-energy solutions are more stable than fixed point solutions, that are not global solutions of the GCP (this is true even without adding noise, although the difference is smaller), under the application of noise.
(a) Average and standard deviation of $O\big{[}\{\tilde{\mathbf{x}}_{i}\}_{i\in\mathcal{V}},\{\mathbf{x}_{i}\}_{i\in\mathcal{V}}\big{]}$ as a function of $\alpha$.
(b) Average and standard deviation of $\Delta O\big{[}\{\mathbf{x}^{FP}_{i}\}_{i\in\mathcal{V}},\{\tilde{\mathbf{x}}^{FP}_{i}\}_{i\in\mathcal{V}},\{\boldsymbol{\xi}_{i}\}_{i\in\mathcal{V}},\{\tilde{\boldsymbol{\xi}}_{i}\}_{i\in\mathcal{V}}\big{]}$ as a function of $\alpha$.
Figure 7: Comparison between the overlap before and after the application of the trained model and the difference of overlaps between zero-energy solution and fixed point solution.
Appendix B Comparison with simulated annealing by iteration
In figure 8 we report the plots showing the scaling of energy with number of iterations for the missing connectivities.
(a) Comparison at connectivity 11.0
(b) Comparison at connectivity 12.0
(c) Comparison at connectivity 12.5
(d) Comparison at connectivity 14.0
(e) Comparison at connectivity 14.5
(f) Comparison at connectivity 15.0
Figure 8: Comparison of the scaling of energy with the number of iterations
Appendix C Comparison with simulated annealing by time
Our algorithm has been executed on a NVIDIA Tesla V100 SXM2 32GB CPU while SA has been executed on a Intel(R) Xeon(R) CPU E5-2698 v4 @ 2.20GHz. In table 2 we report the average number of iterations per second obtained in different configurations for our algorithm and for SA (in parentheses).
Number of nodes Connectivity 11.5 Connectivity 13.0 Connectivity 13.5 Connectivity 15.5
1k 173 231 229 227
3k 211 217 219 218
10k 172 (469) 212 (441) 215 (429) 185 (384)
30k 132 120 117 104
100k 37 33 32 28
Table 2: Average number of iterations per second for different configurations using our method. Simulated annealing times are reported in parentheses.
In figure 9 we report the average time per iteration to execute SA and our method on our hardware.
(a) Comparison at connectivity 11.5
(b) Comparison at connectivity 13.0
(c) Comparison at connectivity 13.5
(d) Comparison at connectivity 15.5
Figure 9: Comparison of the scaling of energy with execution time